20220421_DX_Day. DX AdSec Flattening with New 4D, Day 6

9:19 Getting started

9:20 Changing configuration to daytime one
  • re-introducing actuators: 4, 218, 260, 340, 367, 470, 565
  • changed ff matrix acquisition configuration to 5 iteration of modal with ff

9:25 Restarting ASM to load the 'daytime' configuration and send control to Xianyu

9:26 Enable RR mode for future use, we did not enable_lab_mode and did not stop the AdsecArbitrator

9:27 Acquiring FF with 5 iterations, modal with FF (test)
  • set_plot, 'Z' (to redirect graphical output to /null)
  • print, fsm_set_flat(/no_flat, /ff)
  • tn=tracknum() (20220421_163009)
  • print,ff_acquire_v2(tn)

9:51 Reducing FF data
  • print, reduce_ff_data('20220421_163009', /DO_PLOT)
  • Normalized Chi^2 ff_matrix: 122.57134 DOF: 1602756
  • Maximum FF residual 1.47e-3

9:57 Acquiring FF with 20 iterations, modal with FF
  • tn=tracknum() (20220421_165531)
  • print,ff_acquire_v2(tn)
  • 10:23 Reference Body temperature 5.4 C, Cold Plate temperature 5.0 C
  • 10:41 Reference Body temperature 5.5 C, Cold Plate temperature 5.1 C
  • 11:01 Reference Body temperature 5.6 C, Cold Plate temperature 5.2 C

11:11 Completed acquisition
  • print, fms_rip()

11:12 Reducing FF data
  • print, reduce_ff_data('20220421_165531', /DO_PLOT)
  • Normalized Chi^2 ff_matrix: 123.83306 DOF: 7613091

11:25 Restarting ASM to send control to Doug

11:32 Preparing for IFs measurement test (pupil adjustment)
  • Setting shell with default flat
  • We see decent pupil illumination with many fringes (after some RX, RY correction)
  • We need to adjust the pupil position on the 4D CCD
  • Pupil Mask Position: 3.14, 3.50 (diameter 5.88) central 3.17, 3.49 (diameter 0.581)
  • Pupil (Mirror) Position: 3.18, 3.54 (diameter 5.88)
  • Moving tertiary and correcting with hexapod
  • First coordinated move (delta) Tertiary (Tilt) -90 Secondary (Rx, RY) +15, +21
  • Pupil (Mirror) Position: 3.15, 3.54 (diameter 5.88)
  • Second coordinated move (delta) Tertiary (Tilt) -30 Secondary (Rx, RY) +5, +7
  • Pupil (Mirror) Position: 3.14, 3.54 (diameter 5.88)
  • Third coordinated move (delta) Tertiary (Tip) -90 Secondary (Rx, RY) +31, -20
  • Fringes are back but with some coma
  • Pupil (Mirror) Position: 3.14, 3.50 (diameter 5.88)

12:02 Fine alignment to correct tip-tilt, focus and coma
  • Hexalign
  • Secondary: -3.456, 1.8734, -1.6326, -143.293, -272.3743, -0.0009
  • Tertiary: -218.94, 884.02, 0, 56.408

12:10 Flat application (default)
  • /local/aomeas/adsec_calib/flat/20220420_181045/20220419_225059_bump_low_force.sav
  • WF error (waves at 632 nm, tip-tilt, focus, coma, astigmatism, spherical removed) PV=1.01, RMS=0.136 (-0.20, +0.13, minmax forces)
  • 200 modes re-flattening
  • print, opt_int_mat_reduce2(200, track='20220419_210751’)
  • saved flat: /local/aomeas/adsec_calib/flat/20220421_191459/flattening_result.sav
  • WF error (waves at 632 nm, tip-tilt, focus, coma, astigmatism, spherical removed) PV=0.404, RMS=0.089 (-0.20, +0.12, minmax forces)
  • Note: there is very little trefoil before the re-flattening

12:18 Resting and taking a break
  • print, fms_rip()

14:40 Resuming after AOG meeting

14:50 First test of IFs acquisition
  • Set flat with full 632 modes
  • hexalign
  • 4D interferometer set to return phase maps without any mode removed (shape removal none)
  • i4d_collect with 5 modes [0-4]

15:15 First OPD created successfully
  • i4d_collect
  • print, idl_4d_produce('20220421_215650')
  • print, if_redux_new(/modal,'20220421_215650', plot=1, /force)

15:26 Taking some OPD measurement with different number of modes
  • print, fsm_load_shape("/local/aomeas/adsec_calib/flat/Applicable/default.sav")
  • we tried the command with a limit in the print, fsm_load_shape("/local/aomeas/adsec_calib/flat/Applicable/default.sav")
  • AdOpt>print, fsm_load_shape("/local/aomeas/adsec_calib/flat/Applicable/default.sav",FORCE_MAX_MODE=200)
  • Maximum mode index used to apply flat -> # 200
  • STRTRIM: Variable is undefined: LIMIT_ACT. -10062

15:47 Changing actuator configuration to nighttime
  • 4, 260, 565, 367, 340, 470, 218 NOT in use
  • print, fsm_rip()
  • Power off from Adsec Control GUI
  • adsc_stop/adsc_start
  • Power On from Adsec Control GUI
  • LoadProgram from Adsec Control GUI
  • SetFlat from Adsec Control GUI (default)
  • WF error (waves at 632 nm, tip-tilt, focus, coma, astigmatism, spherical removed) PV=0.788, RMS=0.118 (-0.18, 0.12, minmax forces)
  • 16:05 Reference Body temperature 7.7 C, Cold Plate temperature 7.0 C

16:05 Handed over to Telescope

-- GuidoBrusa - 21 Apr 2022
Topic revision: r2 - 21 Apr 2022, GuidoBrusa
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback